OpenStudy (anonymous):

How many ways can a club of 10 students pick a president, vice president, treasurer, and secretary. I missed this lesson and i need to know how to get this answer : P(10,3) = 720 I am totally lost ;P Please please please pleeeeeeeaaaaase help? :(

OpenStudy (anonymous):

ten choices for president times 9 choices for vp times 8 choices for treasurer times 7 choices for secretary = \[10\times 9\times 8\times 7\]

OpenStudy (anonymous):

Alright which would be 5040 .....

OpenStudy (anonymous):

you should be confused 720 is wrong

OpenStudy (anonymous):

I am on my lesson and it says that the answer.. >.<

OpenStudy (anonymous):

it is P(10,4) not P(10,3) as you have 4 positions to fill, not 3!

OpenStudy (anonymous):

Ik!!! and it says 3!!! thats why im so confused >.<

OpenStudy (anonymous):

it says :

OpenStudy (anonymous):

write the problem in P(n,r) notation then use counting permutation formula to get the numerical answer..

OpenStudy (anonymous):

Does that help any? (: